the ages of a man and his son add up to 39. in 3 years the man will be 4 times as old as his son. Find ages of man and son.

Answer by TimothyLamb(4379)

---
x = man
y = son
---
x + y = 39
x + 3 = 4(y + 3)
---
put the system of linear equations into standard form:
---
x + y = 39
x + 3 = 4y + 12
---
x + y = 39
x - 4y = 9
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = man = 33
y = son = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php